OpenStudy (anonymous):

Suppose that the graph of f contains the point (5,2). Find a point that must be on the graph of g(x)=3f(x+7). How do I solve this?

OpenStudy (anonymous):

I know I gotta plug things in but I am not getting the correct answer so I must be doing it wrong.

hero (hero):

Yeah, there's something I don't like about this

OpenStudy (anonymous):

We just started doing these today in class and the teacher confused the crap outta me like usual.

hero (hero):

Are you doing derivatives?

OpenStudy (anonymous):

no?

hero (hero):

What is the topic of study?

OpenStudy (anonymous):

Transformations

hero (hero):

Okay

hero (hero):

Yeah, I have no clue. I'd have to look up some resources to figure it out

OpenStudy (anonymous):

Thanks for trying anyways...

hero (hero):

I don't see how I'm supposed to know what f(12) means

OpenStudy (anonymous):

f(12)? where did u get that?

hero (hero):

g(5) = 3f(5 + 7)

OpenStudy (anonymous):

ok nevermind

OpenStudy (anonymous):

f contains (5,2) that means \[f(5)=2\]

hero (hero):

I know that already. It wants f(5+7)

hero (hero):

Any ideas?

OpenStudy (anonymous):

no it doesn't ask for that. it says "Find a point that must be on the graph of g(x)=3f(x+7)."

hero (hero):

So....what?

OpenStudy (anonymous):

so the only thing you know is that \[f(5)=2\] so if \[x+7=5\] then \[x=-2\] so \[g(-2)=3f(-2+7)=3f(5)=3\times 2=6\] so the only thing you know is that (-2,6) is on the graph of g

hero (hero):

Hmm. I think if I would have kept going with my original idea, I would have gotten that

OpenStudy (anonymous):

That is correct

hero (hero):

Actually, I wouldn't have. I would have gotten 2,6

OpenStudy (anonymous):

Thanks for trying/solving... might need some more help in a sec